What is the maximum number of phases that can be in equilibrium in a ternary system? Explain.
Answer #1

The answer is 4.

Explanation:

From Gibbs' phase rule we know that F = C – P + N,

where F is the number of degrees of freedom, C is the number of components and P is the number of phases in thermodynamic equilibrium

Now, for a ternary system (C = 3) at constant pressure (N = 1), Gibbs phase rule, becomes

P + F = C + N = 3 + 1 = 4

Or,

P = 4 – F

Thus, when F = 0, P will have its maximum value of 4, which means that the maximum number of phases present for this situation is 4.